Assume 2 kg of hydrogen (Item B 28) are initially at 40 atm, 100 K. (a) If the hydrogen changes pressure to 1 atm without a change of temperature, find the changes of enthalpy and entropy. (b) If the hydrogen changes temperature to 35 K with no change of entropy, find the change of pressure and enthalpy.

1
Expert's answer
2021-11-11T09:12:03-0500


a-

Velocity of the ball at maximum height is zero, v=0

"v \n^2\n \u2212u \n^2\n =2aH"

"0\u2212(49) \n^2\n =2\u00d7(\u22129.8)\u00d7H"

H=122.5 m   


b-

v=u+at

0=49−9.8t

t=5s

Total time taken by ball to return to the surface, T=2t=10 s